You are sitting in a plane, in straight horizontal flight at constant speed of 250 \text{m.s}^{-1}

m.s−1

, when you drop your pen from height 0.30 m above your leg. You've just shown that it hits your leg after 0.25 s.

How far does the plane travel in that time?


1
Expert's answer
2021-10-12T12:32:03-0400

"\text{Let's check correctness of the received result of }"

"\text {time of falling of the pen}"

"h =0.3 m ;g = 9.8\frac{m}{s^2};t =0.25"

"h = \frac{gt^2}{2}= \frac{9.8*0.25^2}{2}= 0.3m"

"t\text{ is correct}"

"v_p = 250\frac{m}{s}; t = 0.25s"

"S = v_p*t = 250*0.25=62.5m"


"\text{Answer: } 62.5m"